\begin{thebibliography}{10}

\bibitem{anwar2024foundational}
Usman Anwar, Abulhair Saparov, Javier Rando, Daniel Paleka, Miles Turpin, Peter Hase, Ekdeep~Singh Lubana, Erik Jenner, Stephen Casper, Oliver Sourbut, et~al.
\newblock Foundational challenges in assuring alignment and safety of large language models.
\newblock {\em Transactions on Machine Learning Research}, 2024.

\bibitem{bai2022constitutional}
Yuntao Bai et~al.
\newblock Constitutional ai: Harmlessness from ai feedback, 2022.

\bibitem{banerjee2025safeinfer}
Somnath Banerjee, Sayan Layek, Soham Tripathy, Shanu Kumar, Animesh Mukherjee, and Rima Hazra.
\newblock Safeinfer: Context adaptive decoding time safety alignment for large language models.
\newblock In {\em AAAI}, volume~39, pages 27188--27196, 2025.

\bibitem{brownLargeLanguageMonkeys2024}
Bradley Brown, Jordan Juravsky, Ryan Ehrlich, Ronald Clark, Quoc~V. Le, Christopher R{\'e}, and Azalia Mirhoseini.
\newblock Large language monkeys: Scaling inference compute with repeated sampling.
\newblock {\em arXiv preprint arXiv:2407.21787}, 2024.

\bibitem{chao2024jailbreakbenchopenrobustnessbenchmark}
Patrick Chao, Edoardo Debenedetti, Alexander Robey, Maksym Andriushchenko, Francesco Croce, Vikash Sehwag, Edgar Dobriban, Nicolas Flammarion, George~J. Pappas, Florian Tramer, Hamed Hassani, and Eric Wong.
\newblock Jailbreakbench: An open robustness benchmark for jailbreaking large language models, 2024.

\bibitem{chen2024finding}
Jianhui Chen, Xiaozhi Wang, Zijun Yao, Yushi Bai, Lei Hou, and Juanzi Li.
\newblock Finding safety neurons in large language models.
\newblock {\em arXiv preprint arXiv:2406.14144}, 2024.

\bibitem{dai2024safe}
Josef Dai, Xuehai Pan, Ruiyang Sun, et~al.
\newblock Safe rlhf: Safe reinforcement learning from human feedback.
\newblock In {\em ICLR}, 2024.

\bibitem{fan-etal-2018-hierarchical}
Angela Fan, Mike Lewis, and Yann Dauphin.
\newblock Hierarchical neural story generation.
\newblock In {\em Proceedings of the 56th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers)}, 2018.

\bibitem{ficler2017controlling}
Jessica Ficler and Yoav Goldberg.
\newblock Controlling linguistic style aspects in neural language generation.
\newblock In {\em Proceedings of the Workshop on Stylistic Variation}, 2017.

\bibitem{greveEvolvingNeuralTuring2016}
Rasmus~Boll Greve, Emil~Juul Jacobsen, and Sebastian Risi.
\newblock Evolving neural turing machines for reward-based learning.
\newblock In {\em Proceedings of the Genetic and Evolutionary Computation Conference 2016}, pages 117--124, Denver Colorado USA, 2016. ACM.

\bibitem{hendrycksmath2021}
Dan Hendrycks, Collin Burns, Saurav Kadavath, Akul Arora, Steven Basart, Eric Tang, Dawn Song, and Jacob Steinhardt.
\newblock Measuring mathematical problem solving with the math dataset.
\newblock In {\em Advances in Neural Information Processing Systems}, 2021.

\bibitem{holtzman2020curious}
Ari Holtzman, Jan Buys, Li~Du, Maxwell Forbes, and Yejin Choi.
\newblock The curious case of neural text degeneration.
\newblock In {\em International Conference on Learning Representations}, 2020.

\bibitem{holtzman-etal-2018-learning}
Ari Holtzman, Jan Buys, Maxwell Forbes, Antoine Bosselut, David Golub, and Yejin Choi.
\newblock Learning to write with cooperative discriminators.
\newblock In {\em Proceedings of the 56th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers)}, 2018.

\bibitem{inan2023llama}
Hakan Inan, Kartikeya Upasani, Jianfeng Chi, Rashi Rungta, Krithika Iyer, Yuning Mao, Michael Tontchev, Qing Hu, Brian Fuller, Davide Testuggine, et~al.
\newblock Llama guard: Llm-based input-output safeguard for human-ai conversations.
\newblock {\em arXiv preprint arXiv:2312.06674}, 2023.

\bibitem{ji2023ai}
Jiaming Ji, Tianyi Qiu, Boyuan Chen, Borong Zhang, Hantao Lou, Kaile Wang, Yawen Duan, Zhonghao He, Jiayi Zhou, Zhaowei Zhang, et~al.
\newblock Ai alignment: A comprehensive survey.
\newblock {\em arXiv preprint arXiv:2310.19852}, 2023.

\bibitem{jiang2023mistral}
Albert~Q Jiang, A~Sablayrolles, A~Mensch, C~Bamford, D~Singh Chaplot, Ddl Casas, F~Bressand, G~Lengyel, G~Lample, L~Saulnier, et~al.
\newblock Mistral 7b. arxiv.
\newblock {\em arXiv preprint arXiv:2310.06825}, 10, 2023.

\bibitem{korbak2023pretraining}
Tomasz Korbak, Kejian Shi, Angelica Chen, Rasika Bhalerao, Christopher~L. Buckley, Jason Phang, Samuel~R. Bowman, and Ethan Perez.
\newblock Pretraining language models with human preferences.
\newblock In {\em ICML}, 2023.

\bibitem{liContrastiveDecodingOpenended2023}
Xiang~Lisa Li, Ari Holtzman, Daniel Fried, Percy Liang, Jason Eisner, Tatsunori Hashimoto, Luke Zettlemoyer, and Mike Lewis.
\newblock Contrastive decoding: Open-ended text generation as optimization.
\newblock In {\em Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers)}, 2023.

\bibitem{liu2024lipo}
Tianqi Liu, Zhen Qin, Junru Wu, Jiaming Shen, Misha Khalman, Rishabh Joshi, Yao Zhao, Mohammad Saleh, Simon Baumgartner, Jialu Liu, et~al.
\newblock Lipo: Listwise preference optimization through learning-to-rank.
\newblock {\em arXiv preprint arXiv:2402.01878}, 2024.

\bibitem{liu2023jailbreaking}
Yi~Liu, Gelei Deng, Zhengzi Xu, Yuekang Li, Yaowen Zheng, Ying Zhang, Lida Zhao, Tianwei Zhang, and Yang Liu.
\newblock Jailbreaking chatgpt via prompt engineering: An empirical study, 2023.

\bibitem{mazeika2024harmbench}
Mantas Mazeika, Long Phan, Xuwang Yin, Andy Zou, Zifan Wang, Norman Mu, Elham Sakhaee, Nathaniel Li, Steven Basart, Bo~Li, David Forsyth, and Dan Hendrycks.
\newblock Harmbench: A standardized evaluation framework for automated red teaming and robust refusal.
\newblock In {\em ICML}, 2024.

\bibitem{panda2024llm}
Swetasudha Panda, Naveen~Jafer Nizar, and Michael~L Wick.
\newblock Llm improvement for jailbreak defense: Analysis through the lens of over-refusal.
\newblock In {\em Neurips Safe Generative AI Workshop 2024}, 2024.

\bibitem{purpura2022learning}
Alberto Purpura, Gianmaria Silvello, and Gian~Antonio Susto.
\newblock Learning to rank from relevance judgments distributions.
\newblock {\em Journal of the Association for Information Science and Technology}, 73(9):1236--1252, 2022.

\bibitem{qi2024safety}
Xiangyu Qi, Ashwinee Panda, Kaifeng Lyu, Xiao Ma, Subhrajit Roy, Ahmad Beirami, Prateek Mittal, and Peter Henderson.
\newblock Safety alignment should be made more than just a few tokens deep.
\newblock {\em arXiv preprint arXiv:2406.05946}, 2024.

\bibitem{radford2019language}
Alec Radford, Jeff Wu, Rewon Child, David Luan, Dario Amodei, and Ilya Sutskever.
\newblock Language models are unsupervised multitask learners.
\newblock 2019.

\bibitem{setlurRewardingProgressScaling2024}
Amrith Setlur, Chirag Nagpal, Adam Fisch, Xinyang Geng, Jacob Eisenstein, Rishabh Agarwal, Alekh Agarwal, Jonathan Berant, and Aviral Kumar.
\newblock Rewarding progress: Scaling automated process verifiers for llm reasoning.
\newblock {\em arXiv preprint arXiv:2410.08146}, 2024.

\bibitem{skeanDoesRepresentationMatter2024}
Oscar Skean, Md~Rifat Arefin, Yann LeCun, and Ravid Shwartz-Ziv.
\newblock Does representation matter? exploring intermediate layers in large language models.
\newblock {\em arXiv preprint arXiv:2412.09563}, 2024.

\bibitem{trungReFTReasoningReinforced2024}
Luong Trung, Xinbo Zhang, Zhanming Jie, Peng Sun, Xiaoran Jin, and Hang Li.
\newblock Reft: Reasoning with reinforced fine-tuning.
\newblock In {\em Proceedings of the 62nd Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers)}, 2024.

\bibitem{wangInterpretablePreferencesMultiObjective2024}
Haoxiang Wang, Wei Xiong, Tengyang Xie, Han Zhao, and Tong Zhang.
\newblock Interpretable preferences via multi-objective reward modeling and mixture-of-experts, 2024.

\bibitem{wangSelfConsistencyImprovesChain2023}
Xuezhi Wang, Jason Wei, Dale Schuurmans, Quoc Le, Ed~Chi, Sharan Narang, Aakanksha Chowdhery, and Denny Zhou.
\newblock Self-consistency improves chain of thought reasoning in language models, 2023.

\bibitem{wang2024chain}
Xuezhi Wang and Denny Zhou.
\newblock Chain-of-thought reasoning without prompting.
\newblock {\em arXiv preprint arXiv:2402.10200}, 2024.

\bibitem{wang2024comprehensive}
Zhichao Wang, Bin Bi, Shiva~Kumar Pentyala, Kiran Ramnath, Sougata Chaudhuri, Shubham Mehrotra, Xiang-Bo Mao, Sitaram Asur, et~al.
\newblock A comprehensive survey of llm alignment techniques: Rlhf, rlaif, ppo, dpo and more.
\newblock {\em arXiv preprint arXiv:2407.16216}, 2024.

\bibitem{wei2023jailbroken}
Alexander Wei, Nika Haghtalab, and Jacob Steinhardt.
\newblock Jailbroken: How does llm safety training fail?
\newblock In {\em NeurIPS}, 2023.

\bibitem{wei2024assessing}
Boyi Wei, Kaixuan Huang, Yangsibo Huang, Tinghao Xie, Xiangyu Qi, Mengzhou Xia, Prateek Mittal, Mengdi Wang, and Peter Henderson.
\newblock Assessing the brittleness of safety alignment via pruning and low-rank modifications.
\newblock In {\em ICML}, 2024.

\bibitem{wei2023jailbreak}
Zeming Wei, Yifei Wang, and Yisen Wang.
\newblock Jailbreak and guard aligned language models with only few in-context demonstrations.
\newblock {\em arXiv preprint arXiv:2310.06387}, 2023.

\bibitem{xie2025sorry}
Tinghao Xie, Xiangyu Qi, Yi~Zeng, Yangsibo Huang, Udari~Madhushani Sehwag, Kaixuan Huang, Luxi He, Boyi Wei, Dacheng Li, Ying Sheng, et~al.
\newblock Sorry-bench: Systematically evaluating large language model safety refusal.
\newblock In {\em ICLR}, 2025.

\bibitem{xie2023defending}
Yueqi Xie, Jingwei Yi, Jiawei Shao, Justin Curl, Lingjuan Lyu, Qifeng Chen, Xing Xie, and Fangzhao Wu.
\newblock Defending chatgpt against jailbreak attack via self-reminders.
\newblock {\em Nature Machine Intelligence}, 2023.

\bibitem{xu2024safedecoding}
Zhangchen Xu, Fengqing Jiang, Luyao Niu, Jinyuan Jia, Bill~Yuchen Lin, and Radha Poovendran.
\newblock Safedecoding: Defending against jailbreak attacks via safety-aware decoding.
\newblock In {\em ACL}, pages 5587--5605. Association for Computational Linguistics (ACL), 2024.

\bibitem{xuSafeDecodingDefendingJailbreak2024}
Zhangchen Xu, Fengqing Jiang, Luyao Niu, Jinyuan Jia, Bill~Yuchen Lin, and Radha Poovendran.
\newblock {S}afe{D}ecoding: Defending against jailbreak attacks via safety-aware decoding.
\newblock In {\em Proceedings of the 62nd Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers)}, 2024.

\bibitem{Qwen2.5}
An~Yang, Baosong Yang, Beichen Zhang, Binyuan Hui, Bo~Zheng, Bowen Yu, Chengyuan Li, Dayiheng Liu, Fei Huang, Haoran Wei, et~al.
\newblock Qwen2.5 technical report.
\newblock {\em arXiv preprint arXiv:2412.15115}, 2024.

\bibitem{yudkowsky2016ai}
Eliezer Yudkowsky.
\newblock The ai alignment problem: why it is hard, and where to start.
\newblock {\em Symbolic Systems Distinguished Speaker}, 4(1), 2016.

\bibitem{zhangReSTMCTSLLMSelfTraining2024}
Dan Zhang, Sining Zhoubian, Ziniu Hu, Yisong Yue, Yuxiao Dong, and Jie Tang.
\newblock Rest-mcts*: Llm self-training via process reward guided tree search, 2024.

\bibitem{zhangGeneralPreferenceModeling2024}
Yifan Zhang, Ge~Zhang, Yue Wu, Kangping Xu, and Quanquan Gu.
\newblock General preference modeling with preference representations for aligning language models.
\newblock {\em arXiv preprint arXiv:2410.02197}, 2024.

\bibitem{zhang2024adversarial}
Yihao Zhang, Zeming Wei, Jun Sun, and Meng Sun.
\newblock Adversarial representation engineering: A general model editing framework for large language models.
\newblock {\em arXiv preprint arXiv:2404.13752}, 2024.

\bibitem{zhangAlleviatingHallucinationsLarge2024}
Yue Zhang, Leyang Cui, Wei Bi, and Shuming Shi.
\newblock Alleviating hallucinations of large language models through induced hallucinations.
\newblock {\em arXiv preprint arXiv:2312.15710}, 2024.

\bibitem{zhao2025identifying}
Yiran Zhao, Wenxuan Zhang, Yuxi Xie, Anirudh Goyal, Kenji Kawaguchi, and Michael Shieh.
\newblock Identifying and tuning safety neurons in large language models.
\newblock In {\em ICLR}, 2025.

\bibitem{zheng2024prompt}
Chujie Zheng, Fan Yin, Hao Zhou, Fandong Meng, Jie Zhou, Kai-Wei Chang, Minlie Huang, and Nanyun Peng.
\newblock On prompt-driven safeguarding for large language models.
\newblock In {\em International Conference on Machine Learning}, pages 61593--61613. PMLR, 2024.

\bibitem{zheng2023judging}
Lianmin Zheng, Wei-Lin Chiang, Ying Sheng, Siyuan Zhuang, Zhanghao Wu, Yonghao Zhuang, Zi~Lin, Zhuohan Li, Dacheng Li, Eric Xing, et~al.
\newblock Judging llm-as-a-judge with mt-bench and chatbot arena.
\newblock {\em Advances in Neural Information Processing Systems}, 36:46595--46623, 2023.

\bibitem{zou2023representation}
Andy Zou, Long Phan, Sarah Chen, James Campbell, Phillip Guo, Richard Ren, Alexander Pan, Xuwang Yin, Mantas Mazeika, Ann-Kathrin Dombrowski, et~al.
\newblock Representation engineering: A top-down approach to ai transparency.
\newblock {\em arXiv preprint arXiv:2310.01405}, 2023.

\bibitem{zou2023universal}
Andy Zou, Zifan Wang, Nicholas Carlini, Milad Nasr, J~Zico Kolter, and Matt Fredrikson.
\newblock Universal and transferable adversarial attacks on aligned language models.
\newblock {\em arXiv preprint arXiv:2307.15043}, 2023.

\end{thebibliography}

\documentclass{article}
\pdfoutput=1

   \usepackage[nonatbib,preprint]{neurips_2025}

\usepackage[utf8]{inputenc} 
\usepackage[T1]{fontenc}    
\usepackage[pagebackref,breaklinks,colorlinks,citecolor=blue]{hyperref}
\usepackage{url}            
\usepackage{booktabs}       
\usepackage{amsfonts}       
\usepackage{nicefrac}       
\usepackage{microtype}      
\usepackage{xcolor}         
\usepackage{marvosym} 

\usepackage{amsmath,amssymb} 
\usepackage{graphicx} 
\usepackage{multirow} 
\usepackage{float} 
\usepackage{caption}
\usepackage{enumitem} 
\usepackage{algorithm}
\usepackage[noend]{algpseudocode}

\title{Advancing LLM Safe Alignment with\\ Safety Representation Ranking}

\author{%
Tianqi Du${}^{1}$\thanks{Equal contribution.}\qquad
Zeming Wei${}^{2*}$\qquad
Quan Chen${}^{2*}$\qquad
Chenheng Zhang${}^{1}$\qquad
Yisen Wang${}^{1,3}$\thanks{Corresponding Author: Yisen Wang (yisen.wang@pku.edu.cn).}
\vspace{10pt}
\\
${}^1$State Key Lab of General Artificial Intelligence,\\
School of Intelligence Science and Technology, Peking University
\\
${}^2$School of Mathematical Sciences, Peking University
\\
${}^3$Institute for Artificial Intelligence, Peking University
}

\begin{document}

\maketitle

\begin{abstract}
The rapid advancement of large language models (LLMs) has demonstrated milestone success in a variety of tasks, yet their potential for generating harmful content has raised significant safety concerns. Existing safety evaluation approaches typically operate directly on textual responses, overlooking the rich information embedded in the model’s internal representations. In this paper, we propose Safety Representation Ranking (SRR), a listwise ranking framework that selects safe responses using hidden states from the LLM itself. SRR encodes both instructions and candidate completions using intermediate transformer representations and ranks candidates via a lightweight similarity-based scorer. Our approach directly leverages internal model states and supervision at the list level to capture subtle safety signals. Experiments across multiple benchmarks show that SRR significantly improves robustness to adversarial prompts. Our code will be available upon publication.
\end{abstract}

\section{Introduction}

Recent large language models (LLMs) have achieved remarkable capabilities across a wide range of tasks. However, this power comes with serious safety and alignment concerns~\cite{wang2024comprehensive,ji2023ai,anwar2024foundational}. By default, LLMs have the potential to generate biased, toxic, or harmful content, and adversarial jailbreak prompts can coax an LLM into violating its own content guidelines~\cite{liu2023jailbreaking,wei2023jailbroken,zou2023universal}. These vulnerabilities persist despite extensive alignment efforts during pre-training and post-training phases~\cite{bai2022constitutional,dai2024safe,korbak2023pretraining}. In practice, the potential for harmful outputs and the ability to bypass built-in safeguards raise significant concerns for deploying LLMs in real-world applications.

To mitigate these safety risks, prior work has explored a variety of defense mechanisms. A common strategy is decoding-time intervention, which redirects the decoding logic of the LLM during inference, through token distributions~\cite{xu2024safedecoding,banerjee2025safeinfer} or safe prompts~\cite{xie2023defending,wei2023jailbreak,zheng2024prompt}. For example, SafeDecoding~\cite{xu2024safedecoding} adjusts the token distribution toward safe response distributions during decoding, while in-context defense~\cite{wei2023jailbreak} aligns the generation distributions to safe contexts with demonstrations. Such interventions can introduce a trade-off between safety and fluency: altering the decoding process may \textbf{degrade the model’s natural performance} on benign inputs or increase inference cost. Meanwhile, post-processing-based defenses apply judging LLMs to inspect the harmfulness of LLMs~\cite{inan2023llama,mazeika2024harmbench}. Unfortunately, recent studies have shown that LLM-based safety judges are often overcautious: they \textbf{flag many benign prompts as unsafe} (so-called over-refusal)~\cite{panda2024llm,xie2025sorry}. This unreliability, \textit{i.e.}, high false-positive rates, limits their practical use, as it can render the model unhelpful even on innocuous tasks. 

In this work, we propose an alternative paradigm (which we call Safety Representation Ranking, SRR) for LLM safety that avoids alteration of the base model’s generation logic and unreliable external judges. Our key idea is to generate multiple candidate responses to a given prompt and then rank them by safety using \textbf{the model’s internal representations}. This approach is similar to using a learned reward model to select outputs~\cite{greveEvolvingNeuralTuring2016,brownLargeLanguageMonkeys2024,zhangReSTMCTSLLMSelfTraining2024}, but there exists an important twist: Traditional reward models are trained on the final generated text, often focusing on general measures of quality or alignment. In contrast, our proposed SRR explicitly targets safety by learning directly from the LLM’s latent features. Existing external reward models may miss fine-grained safety cues embedded in the LLM’s state vectors. Moreover, relying solely on an LLM to judge its own outputs can be unreliable and costly. By delving into the model’s internal representation space, SRR can successfully detect subtle safety-critical representations~\cite{wei2024assessing,zou2023representation,zheng2023judging} that an output-only classifier might overlook, and do so with a lightweight ranking step at inference time.

The SRR framework works in two phases. First, we identify safety-sensitive representations through contrastive training. We construct safety contrastive groups: for each prompt, we use examples of both safe and harmful responses. We feed these paired responses through the LLM and extract their internal representations. Because the groups are semantically related but differ in safety, we can train a lightweight model (a single-layer Transformer) to distinguish \emph{safe} vectors from \emph{unsafe} ones. Through this process, SRR learns which features of the LLM’s latent space correlate with safe content. Then, at inference time, we use the learned safety signals to rank candidate responses. In effect, SRR filters among the model’s own outputs without changing how they were produced. Because it operates on the outputs after generation, SRR imposes almost no modification to the LLM’s decoding logic. Its only overhead is the additional cost of scoring a few extra responses with a small model, which is negligible compared to full decoding.

We conduct comprehensive experiments to validate the effectiveness of the SRR model in identifying the safety responses across multiple datasets. Not only can SRR achieve a sufficiently high accuracy in unseen harmful prompts, but it can also generalize well across different safety evaluation datasets, demonstrating its prominent generalization ability in terms of safety ranking. Additionally, we extend our analysis in terms of other alignment perspectives like privacy and fairness, which validates the potential of SRR for diverse alignment considerations and broadens the applications of SRR.

Grounded by these empirical analyses, we characterize the practicality of SRR for serving as a safeguard module in real-world deployments. First, we incorporate SRR into LLM generation to study how it strengthens their robustness against jailbreak attacks. Additionally, we compare the natural performance of SRR with vanilla generation and other defense paradigms. Because SRR only ranks among natural outputs, the quality and correctness on benign queries remain essentially unchanged. Overall, our empirical results suggest that SRR is both a practical and effective module for LLM alignment.

Our contributions in this paper can be summarized as follows:

\begin{enumerate}
    \item We introduce a novel paradigm, Safety Representation Ranking (SRR), which uses LLM internal representations to rank candidate responses by safety, without altering the model’s decoding logic.
    \item We demonstrate that SRR accurately selects safe outputs across diverse safety benchmarks, generalizes to novel prompts, and can be adapted to other alignment perspectives like privacy and fairness.
    \item We show that integrating SRR into LLM inference significantly reduces harmful outputs under attack, with negligible impact on normal task performance.
\end{enumerate}

\section{Related Work}

\textbf{LLM Safe Alignment}. 
The issue of ensuring safe alignment in LLMs has become a longstanding challenge critical to their trustworthy deployment~\cite{anwar2024foundational,ji2023ai,yudkowsky2016ai}. Specifically, LLMs have shown a tendency to generate harmful responses when confronted with malicious requests. While current alignment techniques have improved at mitigating these risks to some extent, they still tend to be superficial and inadequate~\cite{qi2024safety}. Additionally, inference-time defenses can reduce the success rate of these attacks, but they often struggle with a significant drawback of rejecting benign inputs, leading to over-refusal issues. The underlying mechanism of such issues is that these distribution-based or prompt-based defenses commonly change the decoding strategies of LLMs, making their generation distributions favor refusals. Thus, ensuring safe alignment whilst maintaining the generation distribution stands for a viable solution for these risks.

\textbf{Safety Representations of LLMs}. Building on the representation engineering techniques of LLMs~\cite{zou2023representation,zhang2024adversarial}, which examine LLM dynamics through the lens of hidden space with perspective-specific data, recent research has revealed the existence of safety representations within these models~\cite{wei2024assessing,zheng2024prompt}. Specifically, low-dimensional and structured representations emerge in the hidden states of LLMs, which indicate their safety status. When these representations are activated in specific directions, the LLM can successfully recognize and refuse harmful prompts that go against its ethical guidelines. Conversely, when the activations move in the opposite directions, the LLMs fail to reject harmful inputs and display jailbreak behavior. This interesting property has attracted significant research interest aimed at locating and interpreting these representations~\cite{chen2024finding,zhao2025identifying}. Nonetheless, effective methods for leveraging them to enhance the safety of LLMs remain underexplored.

\textbf{Ranking-based LLM generation}.
A variety of rule-based generation methods have been proposed to improve language model performance, including top-$k$ sampling \cite{fan-etal-2018-hierarchical, holtzman-etal-2018-learning}, temperature-based sampling \cite{ficler2017controlling}, and nucleus sampling \cite{holtzman2020curious}. Beyond these, more refined algorithms have been developed to focus on specific tasks.  For example, \cite{wangSelfConsistencyImprovesChain2023, wang2024chain} leverage majority voting to improve Chain-of-Thought (CoT) reasoning. \cite{xuSafeDecodingDefendingJailbreak2024, liContrastiveDecodingOpenended2023,zhangAlleviatingHallucinationsLarge2024} employ carefully designed decoding methods to generate responses that better align with specific requirements in constrained scenarios.
Recent studies \cite{setlurRewardingProgressScaling2024, wangInterpretablePreferencesMultiObjective2024, zhangGeneralPreferenceModeling2024, trungReFTReasoningReinforced2024} train additional reward models, scaled even equivalently to the base models, to perform reranking for specific tasks.
However, these approaches are either rule-based, task-specific, or impose significant computational overhead, inherently limiting their performance potential and application scope. To overcome these limitations, we propose a more general and lightweight ranker to optimize inference-time computation and extend its applicability across diverse tasks.

\section{Methodology}
\label{section:method}

In this section, we propose Safety Representation Ranking (SRR), a listwise learning-to-rank framework for scoring LLM responses by safety. Given an instruction, SRR generates a set of candidate completions and ranks them such that safe responses receive higher scores than unsafe ones. The core idea is to extract internal representations from a frozen base LLM and train a lightweight transformer ranker to assess instruction-response compatibility. Below, we describe the key components of SRR: candidate response generation, ranker architecture, and optimization with a listwise ranking objective. 

\subsection{Candidate Response Generation} To construct candidate lists for training, we sample the base LLM multiple times using stochastic decoding with moderate temperature. This yields a diverse set of $m$ plausible responses $\{{resp}_1,\dots,{resp}_m\}$. for each instruction. We remove duplicates and include both benign and adversarial candidates by injecting jailbreak prompts~\cite{wei2023jailbreak,zou2023universal}. This helps ensure that the candidate pool contains both safe answers and hard negatives (unsafe answers) for training. Each response is labeled with a binary safety tag $y_i \in \{0,1\}$, where $y_i = 1$ indicates a safe response. For training, we construct tuples of the form $(inst, \{resp_i,y_i\}_{i=1}^m)$, where each list includes at least one safe and one unsafe response.

\subsection{Ranker Model Architecture} The core of SRR is a neural ranker that computes a compatibility score between an instruction and each candidate response. We build this ranker as follows:
\begin{itemize}
    \item \textbf{Step 1. Representation extraction:} We use the base LLM as a fixed feature extractor. For each textual input (instruction or response), we run it through the LLM and take the hidden-state vector at a selected layer as its representation. Concretely, let $\mathbf{h}_\text{inst}\in\mathbb{R}^d$ be the hidden vector for the instruction (the state of the last token in the sequence) at the chosen layer, and let $\mathbf{h}_{\text{resp},i}\in\mathbb{R}^d$ be the hidden vector for the $i$-th response. Since the backbone is trained for next-token prediction, the final layers tend to overfit to this specific task. In contrast, intermediate layers typically provide more comprehensive representations of the preceding context, making them better suited for capturing the overall features required for ranking \cite{skeanDoesRepresentationMatter2024}. Therefore, we adopt intermeidiate layers to capture high-quality semantic content.
    \item \textbf{Step 2. Transformer encoder:} We map each high-dimensional LLM vector (typically $d=4096$) to a lower-dimensional space using a shared learned linear projection. This makes the downstream transformer encoder more lightweight and efficient. We concatenate the projected vectors into a sequence:  
    
    \begin{equation}
       [\mathbf{h}_\text{inst}, \mathbf{h}_{\text{resp},1}, \dots, \mathbf{h}_{\text{resp},m}]. 
    \end{equation}
    
    This sequence is then passed through a Transformer encoder (single-layer in our implementation). The Transformer’s self-attention layers let the instruction embedding interact with each response embedding. After passing through the encoder, we obtain output vectors $\mathbf{o}_\text{inst}$ and $\mathbf{o}_{\text{resp},i}$ corresponding to the instruction and each response, respectively. Intuitively, $\mathbf{o}_\text{inst}$ is the contextualized instruction representation (having attended to all responses) and $\mathbf{o}_{\text{resp},i}$ is the $i$th response representation attended to the instruction.
    \item \textbf{Step 3. Similarity computation:} From these encoder outputs we compute a similarity score $s_i$ for each response. We use cosine similarity:
    
    \begin{equation}
        s_i = \cos(\mathbf{o}_\text{inst}, \mathbf{o}_{\text{resp},i}) = 
\dfrac{\mathbf{o}_\text{inst}^\top \mathbf{o}_{\text{resp},i}}{\|\mathbf{o}_\text{inst}\| \|\mathbf{o}_{\text{resp},i}\|}.
    \end{equation}
    
    These scores $s_i\in[-1,1]$ measure the alignment between instruction and responses in the embedding space, which are used as unnormalized logits for ranking, with a temperature scaling parameter $\tau$ applied before softmax to control sharpness.

\end{itemize}

\subsection{Training Objectives and Pipeline
} 
We train the ranker end-to-end (keeping the base LLM frozen) using a listwise ranking loss. For safe/unsafe training, we interpret the similarity scores ${s_i}$ for a list of $m$ candidates as unnormalized logit scores. We compute a softmax probability for each response:
\begin{equation}
\hat{p}_i = \dfrac{\exp(s_i/\tau)}{\sum_{j=1}^m \exp(s_j/\tau)}.
\end{equation}
We also define a ground-truth probability distribution $p^*$ over the list, which places all mass on the safe responses. For instance, if there are $k$ safe responses among the $m$, we set $p^*_i = 1/k$ for each safe response with $y_i=1$ and $0$ for unsafe ones with $y_i=0$. Then we minimize the Kullback–Leibler divergence:
\begin{equation}
\mathbb{D}_{\mathrm{KL}} \left( p^* \,\|\, \hat{p_i} \right)=\sum_{i=1}^mp^*_i\log\dfrac{p^*_i}{\hat{p}_i}
\end{equation}
This loss, as a standard choice~\cite{purpura2022learning, liu2024lipo}, encourages the model to assign high probability to safe candidates. In effect, the ranker is trained so that the instruction and safe responses have higher cosine similarity than instruction-unsafe pairs.

\subsection{Brief Summary}

In conclusion, SRR learns to map instructions and responses into a joint embedding space where safety alignment is captured by similarity. The use of a transformer encoder allows each response to be scored in the context of the instruction and the other candidates. In the inference stage, for any new instruction and its candidate outputs, the ranker can compute similarity scores and produce a safety-based ranking, without further supervision. A pseudo-algorithm is provided in Algorithm \ref{alg:srr}. First, SRR generates diverse candidate responses for each instruction during the training phase (line 2). It then extracts features from the LLM and uses a transformer-based ranker to compute similarity scores between instructions and responses (line 3-6). These scores are normalized via softmax and compared to ground-truth probabilities to compute a listwise loss, which is used to update the ranker (line 7-10). During inference, the algorithm repeats the feature extraction and similarity computation steps to rank responses based on safety.

\begin{algorithm}[t]
\caption{Safety Representation Ranking (SRR)}
\label{alg:srr}
\begin{algorithmic}[1]
\Require Instruction in training data, LLM $f$, response generator $\mathcal{G}$, ranker $g_\theta$, temperature $\tau$
\vspace{10pt}
\Statex \textbf{Training Phase:}

\For {each instruction in training data}
    \State $\{\text{resp}_1, \dots, \text{resp}_m\} \gets \mathcal{G}(\text{instruction})$ \Comment{Generate diverse candidate responses}
    \State $y_i \gets$ safety label for each $\text{resp}_i$ \Comment{1 for safe, 0 for unsafe}
    \State $\mathbf{h}_\text{inst} \gets f(\text{inst})$, $\mathbf{h}_{\text{resp},i} \gets f(\text{resp}_i)$ for $i = 1 \dots m$ \Comment{Extract LLM features}
    \State $[\mathbf{o}_\text{inst}, \mathbf{o}_{\text{resp},1}, \dots] \gets g_\theta([\mathbf{h}_\text{inst}, \mathbf{h}_{\text{resp},1}, \dots])$ \Comment{Transformer-based contextual encoding}
    \State $s_i \gets \cos(\mathbf{o}_\text{inst}, \mathbf{o}_{\text{resp},i})$ \Comment{Compute cosine similarity score}
    \State $\hat{p}_i \gets \dfrac{\exp(s_i / \tau)}{\sum_j \exp(s_j / \tau)}$ \Comment{Normalize scores via softmax}
    \State $p^*_i \gets \dfrac{1}{k}$ if $y_i = 1$, else $0$ \Comment{Uniform probability on $k$ safe responses}
    \State $\mathcal{L} \gets \text{KL}(p^* \| \hat{p})$ \Comment{Listwise loss}
    \State Update $\theta$ to minimize $\mathcal{L}$
\EndFor
\vspace{10pt}
\Statex \textbf{Inference Phase:}
\State Given a new instruction and candidate $\{\text{resp}_1, \dots, \text{resp}_m\}$
\State Repeat steps 2-6 to compute $s_i$
\State \Return Responses ranked by descending $s_i$
\end{algorithmic}
\end{algorithm}

\section{Evaluation}

In this section, we conduct comprehensive evaluations to show the effectiveness of SRR, including safety, privacy and bias, starting with the overall setup. To further demonstrate the generality of our approach, we also evaluate its ability on other datasets. We also state that the natural performance in math and coding does not deteriorate after attaching the ranker to the model.

\subsection{Experiment Set-up}

\textbf{Models and datasets.} 
In our experiment, we apply three popular LLMs, including (1) Qwen2.5-7b-Instruct~\cite{Qwen2.5}
(2) Mistral-7-v0.3~\cite{jiang2023mistral}, and (3) Vicuna-7b-v1.5~\cite{zheng2023judging}.

For datasets, we apply Harmbench~\cite{mazeika2024harmbench}, SorryBench~\cite{xie2025sorry}, and JailbreakBench~\cite{chao2024jailbreakbenchopenrobustnessbenchmark}.
The HarmBench dataset here refers to the standard section of the Harmbench dataset, which includes 200 different harmful prompts in various areas. The SorryBench and the JailbreakBench have similar contents. For each dataset, we extract 50 of them as the training dataset, and the rest is used as the testing dataset. For each prompt in a dataset, we sample answers from the base model using In-context Attack and In-context Defense, each 20 times. After sampling, we check whether the answers are valid by examining the keywords in them. For safe answers, we examine "Sorry", "unable", "illegal", and "understand". For harmful answers, we examine "sure", "certainly". We then filter the answers using the above criterion to get high-quality data.

\textbf{Metrics}. In the experiments in this section, we request the model to choose between the safe answer and the harmful answer, and verify its correctness, referring to its label.

\textbf{Ranker Settings}. In all experiments, the rankers are implemented using a single Transformer block. The trainable parameters of the ranker model is less than 5M. They operate on features extracted from approximately the bottom 25\% of the base model's layers. During training and evaluation, every data group includes two candidate answers. The ranker is trained to distinguish the answers as safe and harmful. The hyperparameters are set as follows: learning rate is set to 0.001, weight decay is set to 0.0001, dropout is set to 0.1, and momentum is set to 1.0.

\textbf{Baseline}. The baseline of the experiment adopts a reward model to rate answers generated by the base model. A pretrained GPT2 \cite{radford2019language} is used as the reward model in the experiment. Small as it seems, a GPT2 model is still 20 times larger than the ranker model.

\subsection{Overall evaluation}

\begin{table}[t]
    \centering
    \caption{Ranking accuracy of SRR in distinguishing safe and harmful prompts.}
    \begin{tabular}{cc|ccc|c}
    \toprule

     & & \multicolumn{3}{c|}{Model} \\
    Source Dataset & Method &  
    Qwen & Mistral & Vicuna 
    & \textbf{Average} \\
    \midrule
    \multirow{2}{*}{Harmbench} & Baseline & 41.18 & 35.21 & 57.60 & 44.66\\
    & Ours & \textbf{82.35} & \textbf{91.55} & \textbf{90.40} & \textbf{88.10}\\
    \midrule
    
    \multirow{2}{*}{SorryBench} & Baseline & 56.72 & 52.82 & 55.26 & 54.93\\
    & Ours & \textbf{85.57} & \textbf{90.15} & \textbf{87.98} & \textbf{87.90}\\
    \midrule
     \multirow{2}{*}{JailbreakBench} & Baseline &70.00 & 67.39 & 50.00 & 62.46\\
     & Ours & \textbf{80.00} & \textbf{95.65} & \textbf{95.24} & \textbf{90.30}\\
    \bottomrule
    \end{tabular}
    \label{tab:overall}
\end{table}

We use the transformer-architectured ranker to improve the safety of different models on different datasets. As depicted in the \ref{tab:overall}, our method greatly outperforms the reward model in all base models and all datasets. The accuracy of many experiments reach 90\%. Our lightweight method significantly out performs the reward model (gpt2), despite being far smaller in scale. Specifically, when Qwen is used as the base model, the ranker reaches 82.35\%, 91.55\%, 90.40\% respectively on three datasets. Similarly, the results are 85.57\%, 90.15\%, 87.98\% when the base model is Mistral. Finally the performance is 80.00\%, 95.65\%, 95.24\% when the base model is Vicuna.
This implies that rankers can adapt to even larger models.

\subsection{Cross dataset validation}
To further evaluate the generalization capability of our SRR framework across different safety benchmarks, we conduct cross-dataset validation experiments. We apply the ranker trained on one dataset to other unseen datasets. This experimental setup helps us demonstrate whether the model can effectively identify and prioritize safe responses regardless of the dataset's specific characteristics or the types of adversarial prompts it contains.

\begin{table}[t]
    \centering
    \caption{Cross-dataset ranking accuracy of SRR in distinguishing safe and harmful prompts.}
    \begin{tabular}{cc|ccc|c}
    \toprule

     & & \multicolumn{3}{c|}{Model} \\
    Source Dataset & Evaluation Dataset &  
    Qwen & Mistral & Vicuna 
    & \textbf{Average} \\
    \midrule
    \multirow{2}{*}{Harmbench} & SorryBench & 76.96 & 88.06 & 66.04 & 77.02\\
    & JailbreakBench & 80.00 & 93.48 & 85.71 & 86.40\\
    \midrule
    
    \multirow{2}{*}{SorryBench} & Harmbench & 76.47 & 90.14 & 80.00 & 82.20\\
    & JailbreakBench & 77.78 & 89.13 & 76.19 & 81.03\\
    \midrule
     \multirow{2}{*}{JailbreakBench} & HarmBench &79.41 & 89.44 & 90.40 & 86.42\\
     & SorryBench & 72.41 & 87.16 & 78.59 & 79.39\\

    \bottomrule
    \end{tabular}
    \label{tab:cross}
\end{table}

The results in Table \ref{tab:cross} show that our SRR framework achieves consistently strong cross-dataset performance across all three LLMs (Qwen, Mistral, and Vicuna). When trained on one dataset and evaluated on another, SRR maintains a high level of accuracy in distinguishing safe from harmful responses. For instance, a ranker trained on Harmbench achieves 77.02\% average accuracy on SorryBench and 86.40\% on JailbreakBench. Similarly, a ranker trained on SorryBench achieves 82.20\% on Harmbench and 81.03\% on JailbreakBench. This cross-dataset effectiveness demonstrates that SRR's safety signal is not overly specialized to any particular dataset but instead captures generalizable features of safety within the LLM's internal representations.

This ability to generalize across different safety benchmarks is crucial for real-world deployment. In practical applications, LLMs may encounter a wide variety of adversarial prompts that differ significantly from those seen during training. The strong cross-dataset performance of SRR suggests that it can serve as a robust safeguard module, effectively filtering out harmful responses even when the specific types of attacks vary. This provides evidence that SRR's approach of leveraging internal model representations for safety ranking is both versatile and adaptable to diverse safety challenges.

\subsection{Extension to other alignment perspectives}

In this part, we also extend the application of our Safety Representation Ranking (SRR) framework to other critical alignment perspectives beyond general safety, namely privacy and fairness. These dimensions are essential for ensuring that LLMs not only avoid harmful content but also respect user privacy and produce unbiased, equitable responses. Evaluating SRR's effectiveness in these areas helps to demonstrate its versatility and potential for broader alignment applications.

\textbf{Privacy}. To evaluate the potential of SRR in addressing privacy concerns, we conducted experiments on the Harmcopy dataset, which contains prompts related to privacy infringement. The results are presented in Table \ref{tab:privacy}.

\begin{table}[t]
    \centering
    \caption{Ranking accuracy of SRR in distinguishing infringement and benign prompts.}
    \begin{tabular}{c|ccc|c}
    \toprule
     & \multicolumn{3}{c|}{Model} &  \\
    Dataset & Qwen & Mistral & Vicuna & \textbf{Average} \\
    \midrule
    Harmcopy & 98.08 & 95.83 & 89.74 & 94.28 \\
    \bottomrule
    \end{tabular}
    \label{tab:privacy}
\end{table}

The results show that SRR achieves a high accuracy rate in distinguishing between privacy-infringing and benign prompts across all models. In particular, Qwen demonstrates the highest accuracy of 98.08\%, followed by Mistral with 95.83\% and Vicuna with 89.74\%. The average accuracy across all models is 94.28\%, indicating that SRR is effective in identifying privacy-related safety concerns.

This strong performance in the privacy context further validates the generalizability of our approach. The ability to adapt to privacy-specific prompts shows that SRR can capture fine-grained safety signals related to different alignment perspectives beyond just general harmful content. By leveraging the internal representations of LLMs, SRR can effectively identify privacy risks without requiring extensive retraining or modification of the underlying model architecture. This makes it a versatile and efficient solution for enhancing the privacy safeguards in LLM applications.

\textbf{Fairness}.
To assess the effectiveness of SRR in ensuring fairness, we conducted experiments on the BBQ dataset. This dataset is designed to evaluate the model's ability to avoid generating responses that may contain biases or unfair content. The results are presented in Table \ref{tab:fairness}.

\begin{table}[t]
    \centering
    \caption{Ranking accuracy of SRR in distinguishing safe and harmful prompts.}
    \begin{tabular}{c|ccc|c}
    \toprule
     & \multicolumn{3}{c|}{Model} &  \\
    Dataset & Qwen & Mistral & Vicuna & \textbf{Average} \\
    \midrule
    Biasedbenchmark for QA & 54.82 & 52.09 & 50.64 & 52.52 \\
    \bottomrule
    \end{tabular}
    \label{tab:fairness}
\end{table}

The results indicate that SRR achieves moderate accuracy in identifying and mitigating biased or unfair responses. The average accuracy across all models is 52.52\%, which is relatively lower compared to the results obtained in privacy and safety evaluations. This suggests that while SRR demonstrates some capability in detecting fairness-related issues, there is still room for improvement in this area.
This performance highlights the complexity of fairness as a multifaceted alignment perspective. Fairness evaluation involves nuanced considerations of various social contexts, cultural factors, and potential biases that may not be as directly reflected in the internal representations of LLMs as other forms of harmful content. Despite this, SRR shows a foundational ability to distinguish between more and less fair responses, indicating that it can serve as a starting point for more specialized fairness enhancements in LLM applications.

Overall, the results demonstrate the initial potential of SRR in addressing privacy and fairness concerns for future advancement in this critical area of LLM alignment.

\subsection{Brief Summary}

In this section, we have comprehensively evaluated the effectiveness of our proposed SRR framework across various dimensions of LLM safety and alignment. Our experiments demonstrate that SRR achieves significant improvements in identifying and prioritizing safe responses over harmful ones, with high accuracy across multiple safety benchmarks. The cross-dataset validation further confirms the generalizability of SRR, showing its ability to adapt to different types of adversarial prompts and datasets without being overly specialized. Additionally, our extension to privacy-related prompts reveals SRR's potential in mitigating privacy-infringing outputs, achieving a strong accuracy rate. Even in the context of fairness, where the task is more nuanced, SRR shows a foundational capability to distinguish between biased and unbiased responses, though with moderate accuracy that suggests room for further enhancement. Overall, these results highlight SRR's versatility and effectiveness as a safeguard module that can be integrated into LLM inference to significantly reduce harmful outputs under attacks.

\section{Discussion}

This section further discusses the considerations for SRR in practical deployment. We focus on two fundamental problems: 
\begin{enumerate}
    \item To what extent can SRR mitigate safety alignment issues?
    \item How does SRR impact the natural performance of LLMs?
\end{enumerate}

\subsection{Real-world application}

Recall that we mainly apply the classification accuracy as the main metric to evaluate the precision of SRR in ranking the safety of multiple responses. In this part, we further explore how SRR can improve the safety alignment of LLMs, since aligned LLMs have already exhibited certain robustness against harmful prompts. To this end, we incorporate SRR during real-time inference of the protected LLMs, rather than classifying simulated harmful or safe responses. We also consider practical jailbreak attacks to demonstrate the robustness of SRR. The baseline in this experiment is "first accuracy", which means choosing the answer with the highest possibility generated by the base model. The results shown in Table \ref{tab:real-harm}, \ref{tab:real-jail}, and \ref{tab:real-sorry} demonstrate that SRR significantly enhances the safety alignment of LLMs in real-world applications. When integrated into the inference process of protected LLMs, SRR demonstrates robust performance against practical jailbreak attacks. This indicates that SRR can effectively improve the safety mechanisms of LLMs, reducing their vulnerability to adversarial prompts. By leveraging the model's internal representations, SRR provides an efficient and effective safeguard without compromising the natural performance of the LLMs. Overall, these findings support the practical utility of SRR as a valuable tool for improving the safety and reliability of LLMs in real-world scenarios.

\begin{table}[t]
    \centering
    \caption{Real-world ranking accuracy of SRR in distinguishing safe and harmful prompts in HarmBench}
    \begin{tabular}{c|cc|c}
    \toprule
     & \multicolumn{2}{c|}{Model} &  \\
    Method & Qwen & Mistral & \textbf{Average} \\
    \midrule
    First &  82.52 & 54.43 & 68.48\\
    Ranker & \textbf{83.22} & \textbf{63.29} & \textbf{73.26}\\
    \bottomrule
    \end{tabular}
    \label{tab:real-harm}
\end{table}

\begin{table}[t]
    \centering
    \caption{Real-world ranking accuracy of SRR in distinguishing safe and harmful prompts in JailbreakingBench}
    \begin{tabular}{c|cc|c}
    \toprule
     & \multicolumn{2}{c|}{Model} &  \\
    Method & Qwen & Mistral & \textbf{Average} \\
    \midrule
    First &  16.25 & 32.91 & 24.58\\
    Ranker & \textbf{38.75} & \textbf{39.24} & \textbf{39.00}\\
    \bottomrule
    \end{tabular}
    \label{tab:real-jail}
\end{table}

\begin{table}[t]
    \centering
    \caption{Real-world ranking accuracy of SRR in distinguishing safe and harmful prompts in SorryBench}
    \begin{tabular}{c|cc|c}
    \toprule
     & \multicolumn{2}{c|}{Model} &  \\
    Method & Qwen & Mistral & \textbf{Average} \\
    \midrule
    First &  84.28 & 46.22 & 65.25\\
    Ranker & \textbf{86.16} & \textbf{67.23} & \textbf{76.70}\\
    \bottomrule
    \end{tabular}
    \label{tab:real-sorry}
\end{table}

\subsection{Natural performance}
As discussed in earlier sections, a key advantage of SRR is that it does not intervene in the decoding process of the base language model. This allows SRR to be seamlessly applied at inference time without modifying generation behavior, thereby preserving the model's natural task performance. In this section, we empirically validate this claim using a mathematical reasoning benchmark. We evaluate SRR using the MATH dataset \cite{hendrycksmath2021}, which contains 12,500 competition-level math problems spanning seven topics and five difficulty levels. To assess performance, we extract the final answer from each model-generated response and compare it against the ground-truth answer.

We use Qwen2.5-7B-Instruct as the base model. For each instruction, we sample 10 completions and apply the SRR ranker, which is trained solely on safety datasets, to rank them by their predicted safety. The top-ranked response is selected as the final answer. We then compare the answer accuracy of the ranked responses against the accuracy obtained by the base model's default outputs.

The results are shown in Table \ref{tab:math}. Across all settings, the accuracy of the SRR-ranked completions remains nearly identical to the base model's natural accuracy (68.7\%). In fact, slight fluctuations ($\pm$0.2\%) are observed depending on which safety dataset the ranker was trained on, but these differences fall within the margin of noise and do not indicate degradation in performance. Notably, this result holds despite the SRR ranker being trained exclusively on safety supervision signals, without any exposure to mathematical reasoning data. This demonstrates that the SRR scoring mechanism does not introduce unintended bias toward specific task domains or alter the correctness of model outputs in benign settings.

\begin{table}[t]
    \centering
    \caption{Accuracy (\%) on the MATH dataset when responses are ranked using SRR trained on different safety datasets.}
    \begin{tabular}{c|cccc}
    \toprule
    Source Dataset  & Natural & HarmBench & SorryBench & JailbreakBench \\
    \midrule
    Accuracy & 68.7 & 69.1 & 68.5 & 68.6\\
    \bottomrule
    \end{tabular}
    \label{tab:math}
\end{table}

\section{Limitations}
Although SRR performs excellently in enhancing the safety of LLMs, there are still a few minor limitations. SRR might need task-specific fine-tuning for optimal performance in certain situations, although the training cost is low. While it generalizes well across multiple safety benchmark datasets, its adaptability to special-domain safety scenarios requires further testing. Also, SRR's effectiveness partly relies on the LLM generating diverse candidate responses; if the responses lack diversity, SRR's performance may be somewhat affected. Despite these minor limitations, SRR remains a robust and practical solution for boosting LLM safety and reliability in various real-world applications.

\section{Conclusion}

In this paper, we introduced Safety Representation Ranking (SRR), a novel listwise ranking framework that leverages the internal representations of LLMs to select safe responses without altering the model’s decoding logic. Through contrastive training, SRR identifies safety-sensitive features within the LLM’s hidden states and uses them to rank candidate responses based on safety. Our method not only improves robustness against adversarial prompts but also generalizes well across different safety evaluation datasets. Furthermore, SRR demonstrates potential for addressing other alignment perspectives such as privacy and fairness. Experimental results indicate that SRR significantly reduces harmful outputs under attack while maintaining performance on benign tasks. Overall, SRR serves as a practical and effective safeguard module for LLM alignment, offering a new paradigm for enhancing the safety and reliability of LLMs in real-world applications.

\bibliographystyle{plain}
\bibliography{reference}

\appendix

\end{document}